# Cohomology of modular form connections on complex curves

### v3

@@ Abstract.
We consider the reduction cohomology of modular functions defined on complex curves via generalizations of holomorphic connections. The cohomology is explicitly found in terms of higher genus counterparts of elliptic functions as analytic continuations of solutions for functional equations. Examples of modular functions on various genera are provided.
@@
Cohomology; Complex curves; Modular functions; Elliptic functions; Quasi-Jacobi forms
@@ 1. Introduction
The natural problem of computation of continuous cohomologies for non-commutative structures on manifolds has proven to be a subject of great geometrical interest [9, 23, 37, 16, 19, 47]. As it was demonstrated in [16, 47], the ordinary Gelfand-Fuks cohomology of the Lie algebra of holomorphic vector fields on complex manifolds turns to be not the most effective and general one. For Riemann surfaces, and even for higher dimension complex manifolds, the classical cohomology of vector fields becomes trivial [23]. The Lie algebra of holomorphic vector fields does not always work for cohomology. For example, it is zero for a compact Riemann surface of genus higher than one. In [16] Feigin obtained various results concerning (co)-homology of the Lie algebra cosimplicial objects of holomorphic vector fields $Lie(M)$. Taking into account results in previous approaches, it is desirable also to find a way to enrich cohomological structures. This motivates constructions of more refined cohomology description for non-commutative algebraic structures. In [9], it has been proven that the Gelfand-Fuks cohomology $H^{*}(Vect(M))$ of vector fields on a smooth compact manifold $M$ is isomorphic to the singular cohomology of the space of continuous cross sections of a certain fiber bundle over $M$.
The main aim of this paper is to introduce and compute the reduction cohomology of modular functions satisfying certain conditions related to the general form of reduction formulas [50] on complex curves [15, 6, 21, 1]. Due to the structure of modular forms [14, 7, 18] and reduction relations among them (called Zhu recursion [50] in the vertex algebra literature) used in [48, 31, 20, 45], one can form chain complexes of modular forms depending on $n$ variables that are fine enough to describe the local geometry of complex curves. The reduction formulas allow to relate a modular form depending on $n$ variables with a sum of modular forms depending on $n-1$ variables in a linear way. In the vertex algebra setting, one defines so-called $n$-point correlation functions which depend on $n$ given vertex operators [13] and $n$ ”insertion points” on a given Riemann surface and surface moduli.
In contrast to more geometrical methods, e.g., of ordinary cosimplicial cohomology for Lie algebras [16, 47], the reduction cohomology pays more attention to the analytical and modular structure of elements of chain complex spaces. Computational methods involving reduction formulas (some of them recalled in the last Section [50, 31, 20, 45]) proved their effectiveness in conformal field theory, geometrical descriptions of intertwined modules for Lie algebras, and differential geometry of integrable models.
In Section 2 we give the definition of the reduction cohomology as well as a lemma relating it to the cohomology of generalized connections on $M$. The main proposition explicitly expressing the reduction cohomology in terms of spaces of generalized elliptic functions on $M$ is proven. In Section 3 we provide motivating examples of reduction formulas for various modular functions. Results of this paper are useful for cosimplisial cohomology theory of smooth manifolds, generalizations of the Bott-Segal theorem, and have their consequences in conformal field theory [16, 47], deformation theory, non-commutative geometry, modular forms, and the theory of foliations.
@@ 2. The chain complex and cohomology
In Section 3 we collect motivating examples of meromorphic functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ subject to the general patern of the reduction formulas 2.3. The coefficients $f_{l,k,m}(z_{n+1},\mu)$ and operators $T_{l,k,m}(\mu)$ (see below), that can depend in general on an extra variable $z_{n+1}$ according to the reduction formulas, as well as a set $\mu$ of modular parameters. The explicit dependence $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ on ${\bf z}_{n}$ corresponds to a vertex operator algebra and a genus of a compact complex curve $M$. In each case described in Section 3, functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ have certain automorphic properties proven in the vertex operator algebra picture (cf. corresponding references in Section 3). For any other set of coefficients $f_{l,k,m}(z_{n+1},\mu)$, operators $T_{l,k,m}(\mu)$, and functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$, one defines the corresponding chain complex and cohomology. The construction of the chain complex below has a universal manner since it does not depend on explicit modular properties of functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ when they satisfy the conditions of the next Subsection. Note that the coefficients $f_{l,k,m}(z_{n+1},\mu)$, operators $T_{l,k,m}(\mu)$ are introduced to show the general manner of reduction formulas used to construct coboundary operators, as well as to simplify notations. In some cases we can omit corresponding dependence on certain idices if they are not used in particular coefficients $f_{l,k,m}$ and operators $T_{l,k,m}$.
2.1. Chain complex spaces of $n$-variable modular forms
In this Section we introduce spaces for the chain complex of modular functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ depending on $n$ variables and satisfying certain properties on genus $g$ complex curves [11, 49, 50, 33, 34, 31, 20, 45]. Elements of spaces depend on a set of moduli parameters $\mu\in{\mathcal{M}}$ where we denote by $\mathcal{M}$ a subset of the moduli space of genus $g$ complex curve $M$. Mark $n$ points ${\bf p}_{n}=(p_{1},\ldots,p_{n})$ on a compact complex curve $M$ of genus $g$. Denote by ${\bf z}_{n}=(z_{1},\ldots,z_{n})$ local coordinates around ${\bf p}_{n}\in M$. On a complex curve $M$ of genus $g$, we consider the spaces of modular forms depending on $n$ variables
$C^{n}(\mu)=\left\{\mathcal{Z}\left({\bf z}_{n},\mu\right),n\geq 0\right\},$ (2.1)
that satisfy the reduction formulas. Note that in conformal field theory, modular functions depend not only on a choice of local parameters but also on a choice of states of the theory. Consequences of such dependence are illustrated in Section 3. Given a set of coefficients $f_{l,k,m}\left({\bf z}_{n+1},\mu\right)$ and operators $T_{l,k,m}(\mu)$ satisfying the reduction formulas with a function $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ (see below), the co-boundary operator $\delta^{n}({\bf z}_{n+1})$ on $C^{n}(\mu)$-space is defined accordingly for $\mu$-modular functions (cf. particular examples in Appendix 3, [18, 50, 31, 20, 45]).
For $n\geq 0$, and any $z_{n+1}\in\mathbb{C}$, define $\delta^{n}:C^{n}(\mu){\rightarrow}C^{n+1}(\mu)$, given by
@@ 2.1. Chain complex spaces of $n$-variable modular forms
where $l(g)\geq 0$ is a constant depending on $g$, and the meaning of indices $1\leq k\leq n$, $1\leq l\leq l(g)$, $m\geq 0$ is explained below. The coefficient functions $f_{l,k,m}\left({\bf z}_{n+1},\mu\right)$ have a specific form independent of a vertex operator algebra. For each particular genus $g\geq 0$ of $M$ and a type of modular form defined by the moduli parameter $\mu$, corresponding known operator-valued functions $f_{l,k,m}({\bf z}_{n+1},\mu)T_{l,k,m}(\mu).$ change the $k$-th argument of $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ by changing $\mu$. The reduction formulas have the form:
@@
is satisfied for the coboundary operators (2.2) for spaces $C^{n}(\mu)$. Explicitly, the chain condition (2.4) leads to an infinite $n\geq 0$ set of equations involving functions $f_{l,k,m}\left({\bf z}_{n+1},\mu\right)$ and $\mathcal{Z}\left({\bf z}_{n},\mu\right)$:
@@
The spaces of modular forms satisfying the conditions (2.5) constitute a chain complex
$0\longrightarrow C^{0}\stackrel{{\scriptstyle\delta^{0}}}{{\longrightarrow}}C^{1}\stackrel{{\scriptstyle\delta^{1}}}{{\longrightarrow}}\ldots\stackrel{{\scriptstyle\delta^{n-2}}}{{\longrightarrow}}C^{n-1}\stackrel{{\scriptstyle\delta^{n-1}}}{{\longrightarrow}}C^{n}\longrightarrow\ldots.$ (2.6)
@@
the $n$-th reduction cohomology of $\mu$-modular forms on a complex curve $M$. Note that the reduction cohomology can be defined for all types of modular functions satisfying the reduction formulas (2.3). Operators $T_{l,k,m}(\mu)$, $0\leq l\leq l(g)$, $m\geq 0$, $1\leq k\leq n$, form a set of generators of an infinite-dimensional continual Lie algebra $\mathfrak{g}(\mu)$ endowed with a natural grading indexed $l$, $m$. Indeed, we set the space of functions $\mathcal{Z}({\bf z}_{n},\mu)$ as the base algebra [39, 40, 41, 46] for the continual Lie algebra $\mathfrak{g}(\mu)$, and we take its generators as $X_{k,l,m}\left(\mathcal{Z}\left({\bf z}_{n},\mu\right)\right)=T_{l,k,m}(\mu).\mathcal{Z}\left({\bf z}_{n},\mu\right)$, for $0\leq l\leq l(g)$, $m\geq 0$, $1\leq k\leq n$. Then the commutation relations for operators $T_{l,k,m}(\mu).$ $1\leq k\leq n$ inside $\mathcal{Z}({\bf z}_{n},\mu)$ represent the commutation relations of the continual Lie algebra $\mathfrak{g}(\mu)$. Jacobi identities for $\mathfrak{g}(\mu)$ follow from Jacobi identities of the Lie algebra of operators $T_{l,k,m}(\mu)$. For a given set of functions $f_{l,k,m}\left({\bf z}_{n+1},\mu\right)$, $\mathcal{Z}\left({\bf z}_{n},\mu\right)$, and operators $T_{l,k,m}(\mu)$, the identities (2.5) resulting from equations (2.4) will be discussed in a separate paper.
2.2. The geometrical meaning of reduction formulas and conditions (2.5)
In this Section we show that the reduction formulas have the form of multipoint connections generalizing ordinary holomorphic connections on complex curves [21]. Let us define the notion of a multipoint connection which will be useful for identifying reduction cohomology in Section 2.3. Motivated by the definition of a holomorphic connection for a holomorphic bundle [21] over a smooth complex curve $M$, we introduce the definition of a multiple point connection over $M$.
Let $\mathcal{V}$ be a holomorphic vector bundle on $M$, and $M_{0}\subset M$ be its subdomain. Denote by ${\mathcal{S}\mathcal{V}}$ the space of sections of $\mathcal{V}$. A multi-point connection $\mathcal{G}$ on $\mathcal{V}$ is a $\mathbb{C}$-multi-linear map $\mathcal{G}:M^{n}\to\mathbb{C}$, such that for any holomorphic function $f$, and two sections $\phi(p)$ and $\psi(p^{\prime})$ of $\mathcal{V}$ at points $p$ and $p^{\prime}$ on $M_{0}\subset M$ correspondingly, we have
$\sum\limits_{q,q^{\prime}\in M_{0}\subset M}\mathcal{G}\left(f(\psi(q)).\phi(q^{\prime})\right)=f(\psi(p^{\prime}))\;\mathcal{G}\left(\phi(p)\right)+f(\phi(p))\;\mathcal{G}\left(\psi(p^{\prime})\right),$ (2.8)
where the summation on left hand side is performed over locuses of points $q$, $q^{\prime}$ on $M_{0}$. We denote by ${\mathcal{C}on}_{n}$ the space of connections depending on $n$ points defined over $M$.
Geometrically, for a vector bundle $\mathcal{V}$ defined over $M$, a multi-point connection (2.8) relates two sections $\phi$ and $\psi$ at points $p$ and $p^{\prime}$ with a number of sections on $M_{0}\subset M$. We call
$G(\phi,\psi)=f(\phi(p))\;\mathcal{G}\left(\psi(p^{\prime})\right)+f(\psi(p^{\prime}))\;\mathcal{G}\left(\phi(p)\right)-\sum\limits_{q,q^{\prime}\in M_{0}\subset\mathcal{X}}\mathcal{G}\left(f(\psi(q^{\prime})).\phi(q)\right),$
the form of a connection $\mathcal{G}$ depending on $n$ points. The space of connection forms depending on $n$ points will be denoted by $G^{n}$. Here we prove the following

### v5

# Cohomology of multipoint connections on complex curves

@@ Abstract.
Assuming conformal field theory correlation functions defined on complex curves satisfy the recursion formulas, we express the corresponding cohomology theory via the generalizations of holomorphic connections. The cohomology is explicitly found in terms of higher genus counterparts of elliptic functions as analytic continuations of solutions for functional equations. Explicit examples associated to correlation functions on various genera are provided.
@@
Cohomology; Complex curves; Correlation functions; Elliptic functions
@@ 1. Introduction
In the literature, the question of computation of continuous cohomology of non-commutative structures associated to manifolds attracted always wide geometric interest [10, 24, 38, 17, 20, 50]. The classical Gelfand-Fuks cohomology of the Lie algebra of holomorphic vector fields on complex manifolds does not always reaches its purposes [17, 50]. In cases of higher dimension complex manifolds and Riemann surfaces, the classical cohomology of vector fields vanishes [24]. (Co)-homology of the Lie algebra cosimplicial objects of holomorphic vector fields was discussed in [17]. In [10], it has been proven that Gelfand-Fuks cohomology of vector fields on a smooth compact manifold is isomorphic to the singular cohomology of the space of continuous cross sections of a certain fiber bundle over that manifold. Our motivation to introduce and compute more refined cohomology description for non-commutative algebraic structures, is based on results in previous cohomology approaches.
The main aim of this paper is to formulate and compute the recursion cohomology of correlation functions satisfying certain conditions related to the general form of recursion formulas [53] in terms of geometric multipoint connections on complex curves [16, 7, 22, 1]. We are able to construct chain complexes of correlation functions depending on $n$ variables according with the properties of correlation functions [15, 8, 19] and corresponding Zhu recursion relations [51, 32, 21, 46, 47, 48], That properties allow us to describe the local geometry of complex curves. The recursion formulas allow to relate a correlation function depending on $n$ variables with a sum of correlation functions depending on n1n-1 variables in a linear way. In conformal field theory setting, one defines $n$-point correlation functions which depend on $n$ given vertex operators [14] and $n$ insertion points on a given Riemann surface and surface moduli. While more geometric methods are used in the ordinary cosimplicial cohomology for Lie algebras [17, 50], the formulation of the recursion cohomology in terms of multipoint connections introduced in this paper reflects the modular and analytic properties of the corresponding chain complex spaces. Computational methods based on recursion formulas (some of them recalled in the last Section [53, 32, 21, 46]) turned out very effective in conformal field theory. In Section ?? we give the definition of the recursion cohomology, and provide a lemma relating it to the cohomology of generalized connections on $M$. The main proposition explicitly expressing the recursion cohomology in terms of spaces of generalized elliptic functions on $M$ is proven. In Section ?? we provide motivating examples of recursion formulas for various correlation functions. A formulation of the recursion cohomology in terms of multipoint connections and Proposition ?? have their applications in generalizations of the Bott-Segal theorem [10], computations of characteristic classes in cosimplicial cohomology theory of smooth manifolds [17, 50, 1, 7, 16, 20, 22, 24], in conformal field theory [4, 15, 18, 25, 26, 39, 44], deformation theory, non-commutative geometry, modular forms [2, 5, 6, 8, 9, 12, 13, 19, 23, 27, 28, 29, 30, 31, 34, 35, 37, 43, 52], and the theory of foliations [3]. Taking into account the power of recursion formulas in computaions of correlation functions, Proposition ?? allows to find explicitly the cohomology in terms of combinations of coefficient functions of the corresponding recursion formulas in particular cases mentioned in Section ??.
@@ 2. The chain complex and cohomology
2.1. Chain complex spaces of $n$-variable correlation functions
In this Section we introduce the spaces for a chain complex of correlation functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ depending on $n$ variables and satisfying certain properties on genus $g$ complex curves [12, 52, 53, 34, 35, 32, 21, 46, 47, 48]. Elements of spaces depend on a set of moduli parameters $\mu\in{\mathcal{M}}$ where we denote by $\mathcal{M}$ a subset of the moduli space of genus $g$ complex curve $M$. Mark $n$ points ${\bf p}_{n}=(p_{1},\ldots,p_{n})$ on a compact complex curve $M$ of genus $g$. Denote by ${\bf z}_{n}=(z_{1},\ldots,z_{n})$ local coordinates around ${\bf p}_{n}\in M$. On a complex curve $M$ of genus $g$, we consider the spaces
$C^{n}(\mu)=\left\{\mathcal{Z}\left({\bf z}_{n},\mu\right)\right\},$ (2.1)
of correlation functions depending on $n\geq 0$ variables. We assume that elements of (LABEL:cnmu) satisfy the recursion formulas. Note that in conformal field theory, correlation functions depend not only on a choice of local parameters but also on a choice of states of the theory. Consequences of such dependence are illustrated in Section ??. Recall that a recursion formula expresses an $n+1$-point correlation function $\mathcal{Z}\left({\bf z}_{n+1},\mu\right)$ as a linear combination of $n$-point correlation functions by means of coefficient functions $f_{l,k,m}\left({\bf z}_{n+1},\mu\right)$ and insertion operators $T_{l,k,m}(\mu)$ [53, 32, 21, 8, 46, 47, 48] Given a set of coefficients $f_{l,k,m}\left({\bf z}_{n+1},\mu\right)$ and operators $T_{l,k,m}(\mu)$ satisfying the recursion formulas, the co-boundary operator $\delta^{n}({\bf z}_{n+1})$ is defined on $C^{n}(\mu)$-space as an inverse map increasing the number of variables to $n$ (cf. explicit examples in Appendix ??, [19, 53, 32, 21, 46]). For $n\geq 0$, and any $z_{n+1}\in\mathbb{C}$, define $\delta^{n}:C^{n}(\mu){\rightarrow}C^{n+1}(\mu)$, given by
@@ 2.1. Chain complex spaces of $n$-variable correlation functions
where $l(g)\geq 0$ is a constant depending on $g$, and the meaning of indices $1\leq k\leq n$, $1\leq l\leq l(g)$, $m\geq 0$ is explained below. The coefficient functions $f_{l,k,m}\left({\bf z}_{n+1},\mu\right)$ have a specific form independent of a vertex operator algebra. For each particular genus $g\geq 0$ of $M$ and a type of correlation functions defined by the moduli parameter $\mu$, corresponding known operator-valued functions $f_{l,k,m}({\bf z}_{n+1},\mu)\;T_{l,k,m}(\mu).$ change the $k$-th argument of $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ by changing $\mu$. The recursion formulas have the form:
@@
is satisfied for the coboundary operators (LABEL:poros) for spaces $C^{n}(\mu)$. The chain condition (LABEL:torba) gives an infinite $n\geq 0$ set of equations on functions $f_{l,k,m}\left({\bf z}_{n+1},\mu\right)$ and $\mathcal{Z}\left({\bf z}_{n},\mu\right)$:
@@
The spaces of correlation functions satisfying the conditions (2.5) form the complex
$0\longrightarrow C^{0}\mathrel{{\mathop{\longrightarrow}\limits^{\delta^{0}}}}C^{1}\mathrel{{\mathop{\longrightarrow}\limits^{\delta^{1}}}}\ldots\mathrel{{\mathop{\longrightarrow}\limits^{\delta^{n-2}}}}C^{n-1}\mathrel{{\mathop{\longrightarrow}\limits^{\delta^{n-1}}}}C^{n}\longrightarrow\ldots.$ (2.6)
@@
the $n$-th recursion cohomology of correlation functions on a complex curve $M$. Note that the recursion cohomology can be defined for all types of correlation functions satisfying the recursion formulas (LABEL:reduction). Operators $T_{l,k,m}(\mu)$, $0\leq l\leq l(g)$, $m\geq 0$, $1\leq k\leq n$, form a set of generators of an infinite-dimensional continual Lie algebra $\mathfrak{g}(\mu)$ endowed with a natural grading indexed $l$, $m$. Indeed, we set the space of functions $\mathcal{Z}({\bf z}_{n},\mu)$ as the base algebra [40, 41, 42, 49] for the continual Lie algebra $\mathfrak{g}(\mu)$, and we take its generators as $X_{k,l,m}\left(\mathcal{Z}\left({\bf z}_{n},\mu\right)\right)=T_{l,k,m}(\mu).\mathcal{Z}\left({\bf z}_{n},\mu\right)$, for $0\leq l\leq l(g)$, $m\geq 0$, $1\leq k\leq n$. Then the commutation relations for operators $T_{l,k,m}(\mu).$ $1\leq k\leq n$ inside $\mathcal{Z}({\bf z}_{n},\mu)$ represent the commutation relations of the continual Lie algebra $\mathfrak{g}(\mu)$. Jacobi identities for $\mathfrak{g}(\mu)$ follow from Jacobi identities of the Lie algebra of operators $T_{l,k,m}(\mu)$. For a given set of functions $f_{l,k,m}\left({\bf z}_{n+1},\mu\right)$, $\mathcal{Z}\left({\bf z}_{n},\mu\right)$, and operators $T_{l,k,m}(\mu)$, the identities (2.5) resulting from equations (LABEL:torba) will be discussed in a separate paper.
In Section ?? we collect motivating examples of meromorphic functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ subject to the general patern of the recursion formulas (LABEL:reduction). The coefficients $f_{l,k,m}$ $(z_{n+1},\mu)$ and operators $T_{l,k,m}(\mu)$, that can depend in general on an extra variable $z_{n+1}$ according to the recursion formulas, as well as a set $\mu$ of modular parameters. The explicit dependence $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ on ${\bf z}_{n}$ corresponds to a vertex operator algebra and a genus of a compact complex curve $M$. In each case described in Section ??, functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ have certain automorphic properties proven in conformal field theory picture (cf. corresponding references in Section ??). For any other set of coefficients $f_{l,k,m}(z_{n+1},\mu)$, operators $T_{l,k,m}(\mu)$, and functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$, one defines the corresponding chain complex and cohomology. The construction of the chain complex below has a universal manner since it does not depend on explicit modular properties of functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ when they satisfy the conditions of the next Subsection. Note that the coefficients $f_{l,k,m}(z_{n+1},\mu)$, operators $T_{l,k,m}(\mu)$ are introduced to show the general manner of recursion formulas used to construct coboundary operators, as well as to simplify notations. In some cases we can omit corresponding dependence on certain idices if they are not used in particular coefficients $f_{l,k,m}$ and operators $T_{l,k,m}$.
2.2. Recursion relations as mutipoint connections
Let us now show that the recursion formulas we assume for correlation functions are represented in terms of multipoint connections. The notion of a multipoint connection extends usual holomorphic connections on complex curves [22]. Let $\mathcal{E}$ and $\Omega$ bethe sheaves of holomorphic sections of a holomorphic vector bundle $\mathcal{V}\to\mathcal{X}$ and differentials defined correspondingly over a smooth complex variety $\mathcal{X}$. A holomorphic connection $\nabla$ on $\mathcal{E}$ is a $\mathbb{C}$-linear map $\nabla:E\to E\otimes\Omega$, which satisfy Leibniz formula $\nabla(f\phi)=\nabla f\phi+\phi\otimes dz$, for any holomorphic function $f$. A definition of a multipoint connection on a smooth complex curve $M$ is motivated by the definition of a holomorphic connection. For a holomorphic vector bundle $\mathcal{V}$ on $M$, its subdomain $M_{0}\subset M$, and the space $\mathcal{S}\mathcal{V}$} of $\mathcal{V}$-sections, introduce a multi-point connection $\mathcal{G}$ on $\mathcal{V}$ as a map $\mathcal{G}:M^{n}\to\mathbb{C}$. For any holomorphic function $f$, and two sections $\phi(p)$ and $\psi(p^{\prime})$ of $\mathcal{V}$ at points $p$ and $p^{\prime}$ on $M_{0}\subset M$ correspondingly, we assume that
$f(\psi(p^{\prime}))\;\mathcal{G}\left(\phi(p)\right)+f(\phi(p))\;\mathcal{G}\left(\psi(p^{\prime})\right)=\sum\limits_{q,q^{\prime}\in M_{0}\subset M}\mathcal{G}\left(f(\psi(q)).\phi(q^{\prime})\right),$ (2.8)
where the summation is over locuses of points $q$, $q^{\prime}$ on $M_{0}$. Let us denote by ${\mathcal{C}on}_{n}$ the space of multipoint connections. We call $G(\phi,\psi)=f(\phi(p))\;\mathcal{G}\left(\psi(p^{\prime})\right)+f(\psi(p^{\prime}))\;\mathcal{G}\left(\phi(p)\right)-\sum\limits_{q,q^{\prime}\in M_{0}\subset\mathcal{X}}\mathcal{G}\left(f(\psi(q^{\prime})).\phi(q)\right)$ the form of a connection $\mathcal{G}$. The space of multipoint connection forms is be denoted by $G^{n}$. When the sum over $q$, $q^{\prime}$ in the formulas above vanishes, we call such multipoint connections transversal. Next we have